# Surface decomposition method for sensitivity analysis of first-passage dynamic reliability of linear systems

### v1

@@ Abstract
This work presents a novel surface decomposition method for the sensitivity analysis of first-passage dynamic reliability of linear systems subjected to Gaussian random excitations. The method decomposes the sensitivity of first-passage failure probability into a sum of surface integrals over the constrained component limit-state hypersurfaces. The evaluation of these surface integrals can be accomplished, owing to the availability of closed-form linear expressions of both the component limit-state functions and their sensitivities for linear systems. An importance sampling strategy is introduced to further enhance the efficiency for estimating the sum of these surface integrals. The number of function evaluations required for the reliability sensitivity analysis is typically on the order of $10^{2}$ to $10^{3}$. The approach is particularly advantageous when a large number of design parameters are considered, as the results of function evaluations can be reused across different parameters. Two numerical examples are investigated to demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
For first-passage dynamic reliability problems of linear systems subjected to Gaussian random excitations, several advanced sampling techniques have been developed to achieve extremely efficient and highly accurate estimates of the failure probability au2001first , misraji2020application , katafygiotis2006domain , valdebenito2021failure . The common feature of these sampling techniques lies in that they fully exploit the system linearity and excitation Gaussianity. The efficient importance sampling au2001first constructs its importance density as a weighted sum of the theoretically optimal importance densities associated with the component limit-state functions, in which the weight coefficients are proportional to the component failure probabilities. This idea was further extended in directional sampling, leading to the directional importance sampling misraji2020application . The domain decomposition method katafygiotis2006domain decomposes the first-passage failure probability into a sum of domain integrals that can be evaluated by the direct Monte Carlo simulation, and an importance sampling scheme is then introduced to estimate their total contribution. The multidomain importance sampling valdebenito2021failure also leverages the idea of domain decomposition, where the resulting domain integrals are solved by line sampling, and a similar importance sampling strategy is adopted to estimate their sum.
@@
Reliability sensitivity analysis aims to quantify and understand the influence of small perturbations of system parameters on the probability of system failure wu1994computational , lu2008reliability , papaioannou2018reliability , lacaze2015probability , talebiyan2020sampling , which is essential in risk-informed decision making song2009system , taflanidis2011simulation and reliability-based design optimization youn2004selecting , chun2016structural . Sensitivity analysis of the first-passage dynamic reliability, which represents a challenging branch of reliability sensitivity analysis, has received increasing attention in recent years au2005reliability , chun2015parameter , jensen2015reliability , yang2022structural , xian2023reliability . Au au2005reliability presented an efficient simulation approach designed for dynamic-reliability-based sensitivity analysis involving only one or two parameters by artificially constructing an augmented reliability problem. Chun et al. chun2015parameter developed a sequential compounding method for the parametric sensitivity analysis of general system problems, and then applied this method to a first-passage dynamic reliability problem of a linear system. Jensen et al. jensen2015reliability investigated a subset-simulation-based approach to compute the sensitivities of first-excursion failure probability with respect to the distribution parameters of uncertain design variables. Yang et al. yang2022structural extended the probability density evolution method for evaluation of the first-passage dynamic reliability sensitivities of nonlinear structures with respect to deterministic design parameters. Xian and Su xian2023reliability developed an efficient approach for sensitivity analysis of the first-excursion dynamic reliability of fractionally-damped structures incorporating the explicit time-domain method, adjoint variable method and Poisson process method. Furthermore, building upon the developments confined to linear systems under Gaussian random excitations, the directional importance sampling and domain decomposition method have been extended from dynamic reliability analysis to reliability sensitivity analysis in valdebenito2021sensitivity and misraji2025first , respectively.
@@
In essence, reliability sensitivity analysis entails the evaluation of a high-dimensional integral over the limit-state hypersurface, which is generally challenging to compute papaioannou2018reliability , particularly for first-passage dynamic reliability problems characterized by highly non-smooth and nonlinear limit-state hypersurfaces. Fortunately, in the case of linear systems subjected to Gaussian excitations, the above complex surface integrals can be evaluated with remarkable efficiency. This work presents a novel surface decomposition method to accomplish this purpose. The core concept of this method is to decompose the complex surface integral into a sum of simpler surface integrals over the constrained component limit-state hypersurfaces. The so-called constrained component hypersurface corresponds exactly to the active part of the corresponding component hypersurface that contributes to the system limit-state hypersurface. The effectiveness of this strategy lies in the ease of computing these component surface integrals, as closed-form linear expressions of both the component limit-state functions and their sensitivities are available. An importance sampling scheme is further introduced to estimate the total contribution of these component surface integrals. The sensitivity analyses of the first-passage dynamic reliability problems for an oscillator and a 20-degree-of-freedom shear-type structure are conducted to demonstrate the feasibility of the proposed method.
@@ 3.2 Sensitivity of first-passage dynamic reliability
Considering the system limit-state function shown in Eq.(22), the term $\delta(-G(\boldsymbol{x}))$ in Eq.(23) can be expanded as
$\delta(-G(\boldsymbol{x}))=\sum_{k=1}^{m}\sum_{i=1}^{n}\delta(-g_{ki}(\boldsymbol{x}))\mathbb{I}_{ki}(\boldsymbol{x})$ (24)
where $\mathbb{I}_{ki}(\boldsymbol{x})$ denotes the indicator function that gives 1 if $g_{ki}(\boldsymbol{x})$ attains the minimum among all component limit-state functions, and 0 otherwise. The indicator function can be expressed as
@@
The expansion presented in Eq.(24) decomposes the delta function of the system limit-state function into a sum of the delta functions of the individual component limit-state functions, with the indicator functions serving to partition the domain such that only the component reaching the minimum is active within each subdomain. It is noted that the indicator functions shown in Eq.(25) also provide an effective way to partition the system failure domain in the domain decomposition method for first-passage dynamic reliability analysis.
Differentiating the system limit-state function in Eq.(22) with respect to the design parameter $\theta$, and leveraging the indicator functions defined in Eq.(25), one can obtain the sensitivity of the system limit-state function as
$\frac{\partial G(\boldsymbol{x})}{\partial\theta}=\sum_{k=1}^{m}\sum_{i=1}^{n}\frac{\partial g_{ki}(\boldsymbol{x})}{\partial\theta}\mathbb{I}_{ki}(\boldsymbol{x})$ (26)
@@
Substitution of Eqs.(24) and (26) into Eq.(23) yields the expansion for the sensitivity of system failure probability as follows:
where
@@
Essentially, the indicator function $\mathbb{I}_{ki}(\boldsymbol{x})$ imposes additional constraints, restricting the integration to the portion of the hyperplane where all other component limit-state functions remain in their safe domains.
It should be noted that the constrained hyperplane $g_{ki}(\boldsymbol{x})\mathbb{I}_{ki}(\boldsymbol{x})=0$ corresponds exactly to the portion of the system limit-state hypersurface contributed by the corresponding component limit-state hypersurface. The union of all these constrained hyperplanes, which are mutually exclusive, constitutes the complete system limit-state hypersurface. Consequently, computing the sensitivity of the first-passage dynamic reliability amounts to integrating over a non-smooth system limit-state hypersurface. In this sense, Eqs.(29) and (31) can be interpreted as decomposing this complex surface integral into a set of simpler integrals over the constrained hyperplanes—a strategy referred to as the surface decomposition method. The basic idea of the proposed surface decomposition method is further illustrated in Figure 1 using a 2-dimensional toy problem.
Figure 1: Illustration of surface decomposition method using a 2-dimensional problem. Consider a series system limit-state function $G(\boldsymbol{x})=\min_{k=1}^{4}g_{k}(\boldsymbol{x})$. The four component limit-state functions are $g_{1}(\boldsymbol{x})=12-(-2x_{1}+x_{2})$, $g_{2}(\boldsymbol{x})=18-(-x_{1}+3x_{2})$, $g_{3}(\boldsymbol{x})=36-(6x_{1}+7x_{2})$ and $g_{4}(\boldsymbol{x})=10-(2x_{1}-x_{2})$, respectively. The system failure boundary is composed of four constrained component failure boundaries $g_{k}(\boldsymbol{x})\mathbb{I}_{k}(\boldsymbol{x})=0\;(k=1,2,3,4)$, in which the indicator function $\mathbb{I}_{k}(\boldsymbol{x})$ takes 1 if $g_{j}(\boldsymbol{x})>0,\forall j\neq k$, and 0 otherwise. The sensitivity of series system failure probability can be decomposed into four line integrals over these four constrained component failure boundaries, respectively. These line integrals will become surface integrals over the constrained component failure surfaces in higher-dimensional problems.
Thus far, substituting Eqs.(28) and (32) into Eq.(31), the surface integral $\eta_{ki}$ can be directly estimated with the Monte Carlo simulation as follows:
@@
where $\varphi(\cdot)$ is the standard Gaussian PDF; $\beta_{ki}$ is the reliability index associated with the component limit-state function $g_{ki}(\boldsymbol{x})$, whose closed-form solution is given in Eq.(18); and $\boldsymbol{x}_{j}\;(j=1,2,\dots,N_{ki})$ are samples generated from the PDF $f_{\boldsymbol{X}}(\boldsymbol{x})$ lying on the hyperplane $g_{ki}(\boldsymbol{x})=0$, with $N_{ki}$ being the number of samples.
Direct simulation of sample points on the hyperplane $g_{ki}(\boldsymbol{x})=0$ can be accomplished as follows: i) generate a sample point $\boldsymbol{x}$ based on the PDF $f_{\boldsymbol{X}}(\boldsymbol{x})$; ii) decompose $\boldsymbol{x}$ into a component that aligns with the unit vector $\boldsymbol{u}_{ki}$, $\boldsymbol{x}_{\parallel}=(\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki})\boldsymbol{u}_{ki}$, and a component that is perpendicular to $\boldsymbol{x}_{\parallel}$, $\boldsymbol{x}_{\perp}=\boldsymbol{x}-\boldsymbol{x}_{\parallel}$; iii) construct ${\boldsymbol{x}}^{\prime}=\beta_{ki}\boldsymbol{u}_{ki}+\boldsymbol{x}_{\perp}$, which represents a sample point lying independently on the hyperplane $g_{ki}(\boldsymbol{x})=0$.
@@
The rationale of this importance sampling PMF is that a higher component failure probability $\mathbb{P}_{ki}$ typically indicates a larger measure of the constrained hyperplane $g_{ki}(\boldsymbol{x})\mathbb{I}_{ki}(\boldsymbol{x})=0$. Consequently, the corresponding surface integral $\eta_{ki}$ contributes more significantly to the sensitivity of system failure probability.
@@
where $N$ is the number of samples; $(K_{j},I_{j})$ is the index pair sampled from the importance sampling PMF $h(k,i)$ in Eq.(38); and $\boldsymbol{x}_{j}$ is sampled from the hyperplane $g_{K_{j}I_{j}}(\boldsymbol{x})=0$.

### v2

@@ Abstract
This work presents a novel surface decomposition method for the sensitivity analysis of first-passage dynamic reliability of linear systems subjected to Gaussian random excitations. The method decomposes the sensitivity of first-passage failure probability into a sum of surface integrals over the constrained component limit-state hypersurfaces. The evaluation of these surface integrals can be accomplished, owing to the availability of closed-form linear expressions of both the component limit-state functions and their sensitivities for linear systems. An importance sampling strategy is introduced to further enhance the efficiency for estimating the sum of these surface integrals. The number of function evaluations required for the reliability sensitivity analysis is typically on the order of $10^{2}$ to $10^{3}$. The approach is particularly advantageous when a large number of design parameters are considered, as the results of function evaluations can be reused across different parameters. Three numerical examples are investigated to demonstrate the effectiveness of the proposed method.
@@ 1 Introduction
For first-passage dynamic reliability problems of linear systems subjected to Gaussian random excitations, several advanced sampling techniques have been developed to achieve extremely efficient and highly accurate estimates of the failure probability au2001first, misraji2020application, katafygiotis2006domain, valdebenito2021failure. The common feature of these sampling techniques lies in that they fully exploit the system linearity and excitation Gaussianity. The efficient importance sampling au2001first constructs its importance density as a weighted sum of the theoretically optimal importance densities associated with the component limit-state functions, in which the weight coefficients are proportional to the component failure probabilities. This idea was further extended in directional sampling, leading to the directional importance sampling misraji2020application. The domain decomposition method katafygiotis2006domain decomposes the first-passage failure probability into a sum of domain integrals that can be evaluated by the direct Monte Carlo simulation, and an importance sampling scheme is then introduced to estimate their total contribution. The multidomain line sampling valdebenito2021failure also leverages the idea of domain decomposition, where the resulting domain integrals are solved by line sampling, and a similar importance sampling strategy is adopted to estimate their sum.
Reliability sensitivity analysis aims to quantify and understand the influence of small perturbations of system parameters on the probability of system failure wu1994computational, lacaze2015probability, talebiyan2020sampling, which is essential in risk-informed decision making song2009system, taflanidis2011simulation and reliability-based design optimization youn2004selecting, chun2016structural. Several sampling-based and surrogate-modeling-based methods have been developed to address the reliability sensitivity problems. In the former category, Lu et al. lu2008reliability and Song et al. song2009subset extended line sampling and subset simulation, respectively, for reliability sensitivity analysis with respect to the distribution parameters of basic random variables. Iason et al. papaioannou2018reliability developed the sequential importance sampling for estimating reliability sensitivities with respect to the design parameters that appear in the indicator function, in which a smooth approximation of the indicator function is employed. In the latter category, Rahman rahman2009stochastic investigated the moment and reliability sensitivities with respect to probability distribution parameters leveraging surrogate performance function constructed through dimensional decomposition. Dubourg and Sudret dubourg2014meta incorporated the Kriging-based importance sampling with a score function approach to study the influence of the design parameters on the structural failure probability. Yun et al. yun2020adaptive developed an approach combining adaptive subdomain sampling and adaptive Kriging construction to compute the failure probability sensitivities. Torii et al. torii2017probability approximated both limit-state function and its derivative using polynomial expansions to facilitate sensitivity analysis of the structural failure probability. It is noted that, similar to their application in reliability analysis, the surrogate-modeling-based methods for reliability sensitivity analysis are generally restricted to problems with low-dimensional uncertain input.
@@
Sensitivity analysis of the first-passage dynamic reliability, which represents a challenging branch of reliability sensitivity analysis, has received increasing attention in recent years au2005reliability, jensen2015reliability, chun2015parameter, yang2022structural, li2023direct, xian2023reliability. In the context of sampling-based methods, Au au2005reliability presented an efficient simulation approach designed for dynamic-reliability-based sensitivity analysis involving one or two parameters by artificially constructing an augmented reliability problem. Jensen et al. jensen2015reliability investigated a subset-simulation-based approach to compute the sensitivities of the first-excursion failure probability with respect to the distribution parameters of uncertain design variables. Besides, Chun et al. chun2015parameter developed a sequential compounding method for the parametric sensitivity analysis of general system problems, and then applied this method to a first-passage dynamic reliability problem of a linear system. Yang et al. yang2022structural extended the probability density evolution method for evaluation of the first-passage dynamic reliability sensitivities of nonlinear structures with respect to deterministic design parameters. Li et al. li2023direct also extended the direct probability integral method for first-passage failure probability sensitivity analysis of nonlinear structures. Xian and Su xian2023reliability developed an efficient approach for sensitivity analysis of the first-excursion dynamic reliability of fractionally-damped structures incorporating the explicit time-domain method, adjoint variable method and Poisson process method. Furthermore, building upon the developments confined to linear systems under Gaussian random excitations, the directional importance sampling and domain decomposition method have been extended from dynamic reliability analysis to reliability sensitivity analysis in valdebenito2021sensitivity and misraji2025first, respectively.
@@
In essence, reliability sensitivity analysis entails the evaluation of a high-dimensional integral over the limit-state hypersurface, which is generally challenging to compute papaioannou2018reliability, particularly for first-passage dynamic reliability problems characterized by highly non-smooth and nonlinear limit-state hypersurfaces. Fortunately, in the case of linear systems subjected to Gaussian excitations, the above complex surface integrals can be evaluated with remarkable efficiency. This work presents a novel surface decomposition method to accomplish this purpose. The core concept of this method is to decompose the complex surface integral into a sum of simpler surface integrals over the constrained component limit-state hypersurfaces. The so-called constrained component hypersurface corresponds exactly to the active part of the corresponding component hypersurface that contributes to the system limit-state hypersurface. The effectiveness of this strategy lies in the ease of computing these component surface integrals, as closed-form linear expressions of both the component limit-state functions and their sensitivities are available. An importance sampling scheme is further introduced to estimate the total contribution of these component surface integrals. The sensitivity analyses of the first-passage dynamic reliability problems for an oscillator, a shear-type structure and a building frame structure are conducted to demonstrate the feasibility of the proposed method.
@@ 3.2 Sensitivity of first-passage dynamic reliability
Introduce an indicator function $\mathbb{I}_{ki}(\boldsymbol{x})$, which takes 1 if the corresponding component limit-state function $g_{ki}(\boldsymbol{x})$ attains the minimum among all component limit-state function, and 0 otherwise. The indicator function can be mathematically expressed as
@@
The above indicator functions serve to partition the integration domain of Eq.(23) into $mn$ mutually exclusive and collectively exhaustive subdomains $\Omega_{ki}=\left\{\boldsymbol{x}\in\mathbb{R}^{n}:\mathbb{I}_{ki}(\boldsymbol{x})=1\right\}$ $(i=1,2,\dots,n;k=1,2,\dots,m)$, within each of which only the component attaining the minimum remains active.
With the indicator functions shown in Eq.(24), the sensitivity of system failure probability shown in Eq.(23) can be expanded as
where
$\eta_{ki}=-\int_{\mathbb{R}^{n}}\delta(-G(\boldsymbol{x}))\frac{\partial G(\boldsymbol{x})}{\partial\theta}\mathbb{I}_{ki}(\boldsymbol{x})f_{\boldsymbol{X}}(\boldsymbol{x})\,d\boldsymbol{x}\quad(i=1,2,\dots,n;k=1,2,\dots,m)$ (26)
denotes the $d$-dimensional volume integral over the subdomain $\Omega_{ki}$.
Within each subdomain $\Omega_{ki}$, one has $G(\boldsymbol{x})=g_{ki}(\boldsymbol{x})$, and hence Eq.(26) can be further written as
@@
Essentially, the indicator function $\mathbb{I}_{ki}(\boldsymbol{x})$ imposes additional constraints, restricting the integration to the portion of the hyperplane where all other component limit-state functions remain in their safe domains. Such constrained hyperplanes are denoted as $S_{ki}=\left\{\boldsymbol{x}\in\mathbb{R}^{n}:g_{ki}(\boldsymbol{x})\mathbb{I}_{ki}(\boldsymbol{x})=0\right\}\;(i=1,2,\dots,n;k=1,2,\dots,m)$.
Note that the constrained hyperplane $S_{ki}$ corresponds exactly to the portion of the system limit-state hypersurface contributed by the corresponding component limit-state hypersurface. The union of all constrained hyperplanes $S_{ki}\;(i=1,2,\dots,n;k=1,2,\dots,m)$, which are mutually exclusive, constitutes the complete system limit-state hypersurface. Consequently, computing the sensitivity of the first-passage dynamic reliability amounts to integrating over a non-smooth system limit-state hypersurface. In this sense, Eqs.(25) and (30) can be interpreted as decomposing this complex surface integral into a set of simpler integrals over the constrained hyperplanes—a strategy referred to as the surface decomposition method. The basic idea of the proposed surface decomposition method is further illustrated in Figure 1 using a 3-dimensional toy problem.
Figure 1: Illustration of surface decomposition method using a 3-dimensional problem. Consider a series system limit-state function $G(\boldsymbol{x})=\min_{k=1}^{4}g_{k}(\boldsymbol{x})$. The four component limit-state functions are $g_{1}(\boldsymbol{x})=12-(-2x_{1}+x_{2}+0x_{3})$, $g_{2}(\boldsymbol{x})=18-(-x_{1}+3x_{2}+0x_{3})$, $g_{3}(\boldsymbol{x})=36-(6x_{1}+7x_{2}+0x_{3})$ and $g_{4}(\boldsymbol{x})=10-(2x_{1}-x_{2}+0x_{3})$, respectively. The system failure surface is composed of four constrained component failure planes $g_{k}(\boldsymbol{x})\mathbb{I}_{k}(\boldsymbol{x})=0\;(k=1,2,3,4)$, in which the indicator function $\mathbb{I}_{k}(\boldsymbol{x})$ takes 1 if $g_{j}(\boldsymbol{x})>0,\forall j\neq k$, and 0 otherwise. The sensitivity of series system failure probability can be decomposed into four surface integrals over these four constrained component failure planes, respectively. These component failure planes will become component failure hyperplanes in higher-dimensional problems.
By substituting Eqs.(29) and (31) into Eq.(30), and applying the law of conditional probability, the surface integral $\eta_{ki}$ can be further expressed as
$\begin{split}\eta_{ki}=-\int_{g_{ki}(\boldsymbol{x})=0}\frac{\boldsymbol{b}_{i}^{s_{k}}\boldsymbol{x}}{\left\|\boldsymbol{a}_{i}^{s_{k}}\right\|}\mathbb{I}_{ki}(\boldsymbol{x})f_{\boldsymbol{X}}(\boldsymbol{x}|\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki}=\beta_{ki})\varphi(\beta_{ki})\,dS(\boldsymbol{x})\\ \hfill(i=1,2,\dots,n;\;k=1,2,\dots,m)\end{split}$ (33)
where $\varphi(\cdot)$ is the standard Gaussian PDF; $\beta_{ki}$ and $\boldsymbol{u}_{ki}$ are the reliability index and the unit vector associated with the component limit-state function $g_{ki}(\boldsymbol{x})$, whose closed-form solutions are given in Eqs.(18) and (20), respectively; and $f_{\boldsymbol{X}}(\boldsymbol{x}|\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki}=\beta_{ki})$ is the PDF of $\boldsymbol{X}$ conditional on the fact that it lies on the hyperplane $g_{ki}(\boldsymbol{x})=0$.
Thus far, the surface integral $\eta_{ki}$ in Eq.(33) can be directly estimated with the Monte Carlo simulation as follows:
@@
where $\boldsymbol{x}_{j}\;(j=1,2,\dots,N_{ki})$ denote samples generated from the PDF $f_{\boldsymbol{X}}(\boldsymbol{x}|\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki}=\beta_{ki})$, with $N_{ki}$ being the number of samples.
Simulating sample points from $f_{\boldsymbol{X}}(\boldsymbol{x}|\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki}=\beta_{ki})$ is equivalent to generating sample points from $f_{\boldsymbol{X}}(\boldsymbol{x})$ and then projecting them onto the hyperplane $g_{ki}(\boldsymbol{x})=0$, which can be accomplished as follows: i) generate a sample point $\boldsymbol{x}$ based on the PDF $f_{\boldsymbol{X}}(\boldsymbol{x})$; ii) decompose $\boldsymbol{x}$ into a component that aligns with the unit vector $\boldsymbol{u}_{ki}$, $\boldsymbol{x}_{\parallel}=(\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki})\boldsymbol{u}_{ki}$, and a component that is perpendicular to $\boldsymbol{x}_{\parallel}$, $\boldsymbol{x}_{\perp}=\boldsymbol{x}-\boldsymbol{x}_{\parallel}$; iii) construct ${\boldsymbol{x}}^{\prime}=\beta_{ki}\boldsymbol{u}_{ki}+\boldsymbol{x}_{\perp}$, which represents a sample point independently generated from the conditional PDF $f_{\boldsymbol{X}}(\boldsymbol{x}|\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{ki}=\beta_{ki})$.
@@
The rationale of this importance sampling PMF lies in the fact that a higher component failure probability $\mathbb{P}_{ki}$ typically indicates a larger measure of the constrained hyperplane $S_{ki}$. Consequently, the corresponding surface integral $\eta_{ki}$ contributes more significantly to the sensitivity of system failure probability.
@@
where $N$ is the number of samples; $(K_{j},I_{j})$ is the index pair sampled from the importance sampling PMF $h(k,i)$ in Eq.(38); and $\boldsymbol{x}_{j}$ is sampled from the conditional PDF $f_{\boldsymbol{X}}(\boldsymbol{x}|\boldsymbol{x}^{\mathrm{T}}\boldsymbol{u}_{K_{j}I_{j}}=\beta_{K_{j}I_{j}})$.
@@
It is worth noting that, when a large number of design parameters are considered, an equal number of impulse response sensitivity time-history analyses need to be conducted in Step ii). This initial setup cost can become computationally expensive for large-scale linear systems. To address this issue, the adjoint variable method developed in hu2016explicit can be employed to reduce the computational cost, in which the numerous impulse response sensitivity time-history analyses based on the equation of motion are replaced by a single impulse response time-history analysis governed by an adjoint equation that has the same form as the original motion equation. Therefore, owing to the merit of the adjoint variable method, the initial setup cost of the surface decomposition method is also not associated with the number of design parameters.